### v1

# An FPGA-based Versatile and Digitalized Method for Pulse Laser Repetition Frequency Locking

Abstract
This paper introduces a novel method for locking the repetition frequency of pulse lasers, adaptable to different frequencies, offering significant improvements in system integration and measurement accuracy. The method consists of two primary components: an error amplification module (EAM) and a digital frequency locking module (DFLM) based on FPGA. The EAM integrates a configurable frequency generator (CFG), a configurable frequency multiplier (CFM) and a mixer, to process the laser pulse alongside a high-stability reference source, such as an atomic clock. By employing frequency multiplication and mixing, the EAM amplifies the laser’s frequency error and performs frequency down-conversion, enhancing measurement sensitivity and reducing the hardware requirements of the back-end. The CFG, implemented on a phase-locked loop (PLL) chip, allows for parameter adjustments to accommodate various laser frequencies. The DFLM processes the output from the EAM using a high-speed, ADC-based dual-mixer time-difference (DMTD) method to precisely measure frequency errors. A digital proportional-integral-derivative (PID) controller then provides feedback to achieve accurate frequency locking. To evaluate the proposed method, an FPGA-based electronic system was developed and tested. In laboratory experiment with a custom-built femtosecond fiber laser, the system demonstrated robust locking of the repetition rate, achieving an Allan deviation improvement from $1.51\times 10^{-7}$ to $1.12\times 10^{-12}$ at a gate time of 10 s. Further testing with a voltage-controlled oscillator (VCO) confirmed a long-term stability of $9.58\times 10^{-14}@10s$.
Index Terms:
Field programmable gate array (FPGA), laser frequency locking, frequency error amplifier, digital electronic circuits, dual-mixer time-difference (DMTD).
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Pulse lasers characterized by narrow-lines and high frequency stability are essential components in optical frequency combs. They have widespread applications in both industrial settings and scientific research, including high-precision measurement[1][2], optical communication[3][4], high-resolution optical spectroscopy[5][6], etc. Optical frequency combs are advanced tools that generate millions of mutually coherent and evenly spaced modes, enabling precise phase-coherent connections across the optical, microwave, and terahertz frequency ranges. These frequency combs are essential for a variety of applications and fundamental scientific research[7][8]. However, optic combs have been limited by various factors such as strain, mechanical perturbations[9], temperature and humidity fluctuations[10]. These environmentally induced variations can cause repetition frequency drift in the comb, leading to instability and inaccuracy in measurements[11][12]. The frequency instability, particularly the resulting jitter in the laser, can significantly degrade measurement resolution. Furthermore, absolute distance measurement experiments often need to be performed outdoors, where environmental control is limited[13]. Therefore, a highly integrated, robust and easy-tunable system which can stabilize the laser frequency is very advantageous and necessary in these applications.
Over the past few decades, various methods have been developed to stabilize pulse laser frequency. A common approach is the use of phase-locked loop (PLL) [14][15]. They can effectively lock the frequency of the laser to a specific value. However, cutting-edge experiments are often dynamic and complex, with changing requirements between measurements due to even minor differences in initial input parameters[16]. Traditional PLLs rely on analog electronics to track and correct phase shifts, but they are sensitive to environmental factors, such as cavity temperature fluctuations, making long-term stability difficult to maintain[14]. Additionally, analog components offer limited flexibility for adjustments. To address these limitations, digital PLLs (DPLLs) have been introduced, offering more flexibility and better resilience to environmental disturbances. However, they are still constrained by frequency dividers and multipliers, which hinder high-resolution frequency adjustments.
Currently, digital systems utilizing Micro-Controller Units (MCUs) or FPGAs are preferred for frequency-locked laser systems. These systems leverage their computational power to implement modern frequency measurement and feedback algorithms, achieving precise frequency locking. Digital systems measure frequency and provide feedback through ADCs and DACs, resulting in quantifiable error signals and adjustable feedback parameters that enhance flexibility and accuracy. While MCUs have been successfully employed in laser frequency stabilization using various locking algorithms[17][18], their feedback bandwidth is typically limited to the kHz range[19], restricting their performance in high-bandwidth applications.
To achieve high feedback bandwidth, FPGAs are an ideal platform for high-speed frequency locking circuits, thanks to their parallel processing capabilities and low latency[20]. Locking systems based on FPGAs can reduce system size while also offer high resolution and accuracy in data input and control output, handle high-concurrency data from high-speed ADC, provide real-time feedback processing, and support high-speed data transmission. Furthermore, modern FPGAs are programmable logic devices rich in computational resources, enabling the hardware implementation of complex algorithms required for laser frequency stabilization and easily configured to meet the requirements of different lasers.
The typical FPGA-based platforms like the Red Pitaya STEMlab have been increasingly adopted for real-world applications[21][22], including lock-in amplification, laser frequency stabilization, and intensity stabilization[23][24]. These platforms enable fully digital processing, which minimizes system noise while offering satisfactory timing resolution and high bandwidth[23][24]. However, lasers operate at various repetition frequencies, and fully digital methods face limitations due to lower ADC sampling rates. To prevent spectral aliasing, signals can only be captured below half the ADC’s sampling rate, which reduces adaptability to higher laser repetition frequencies. Simply increasing the ADC sampling rate is impractical, while it would compromise sampling accuracy. To address this issue, common approaches include frequency divider or mixing to lower the input signal frequency[8][10]. However, employing a frequency divider reduces the repetition frequency of the laser while also decreasing the error frequency that needs to be extracted, which leads to a loss in measurement accuracy. On the other hand, the mixing method necessitates the use of signal synthesizers and mixers, which may introduce additional noise.
Figure 1: Schematic of locking circuits for laser frequency stabilization. PD: photodiode, BPF: band-pass filter, LPF: low-pass filter, CFM: configurable frequency multiplier, CFG: configurable frequency generator, PZT: piezoelectric transducer, HV: high voltage, Amp: voltage amplifier, ADC: analog-to-digital converter, DAC: digital-to-analog converter, DMTD: dual-mixer time-difference, PID: proportional integral derivative.
This paper presents an FPGA-based compact laser frequency locking method designed to adapt to various laser repetition frequencies while maintaining high precision and portability. The method utilizes error amplification and the dual-mixer time-difference (DMTD) method. The error amplification module (EAM) multiplies the frequency to amplify the error and subsequently down-converts it, reducing back-end hardware requirements. This design allows the system to adapt to various laser repetition frequencies while significantly enhancing measurement sensitivity with relatively low sampling rates. The method also integrates a digital frequency locking module (DFLM), include two high-speed, high-precision ADCs, a proportional-integral-derivative (PID) controller, and a high-precision DAC for feedback. All data processing algorithms are implemented on the FPGA, ensuring efficient real-time performance. System parameters are adjustable, allowing flexible measurement and frequency locking across a wide range of laser frequencies.
II Methodology
II-A Overall Design
The schematic of the locking circuits for laser frequency stabilization is illustrated in Fig.1. This method measures the frequency difference between the laser repetition frequency and a high-stability reference source, an atomic clock in this case, enabling precise locking of the laser frequency. The pulse of the laser is first converted to electrical signal using a photodiode (PD). This signal is then processed by the EAM, which amplifies and down-converts the frequency error. The configurable frequency multiplier (CFM) performs frequency multiplication on the laser signal, producing high-frequency signals for mixing. The configurable frequency generator (CFG) uses the atomic clock to produce a stable high frequency signal also for mixing. After mixing, a low-pass filter (LPF) is used to obtain the amplified error signal. Through this process, the frequency error is amplified while the frequency of the signal to be sampled is lowered, thus reducing the requirements of the back-end circuitry.
The error signal, alongside the atomic clock signal, is sampled by a high-speed ADC for processing by the FPGA. The FPGA utilizes the DMTD algorithm to calculate the phase difference between the sampled signals. This phase difference is converted into a frequency difference, passed through a low-pass filter, and then fed to the PID module, which controls a high-speed, high-precision DAC for feedback. Depending on the laser’s repetition frequency and the feedback devices employed, CFM, CFG and digital units within the FPGA can be configured to accommodate them. This capability facilitates the configuration of the EAM and the digital units within the FPGA, ensuring the method’s adaptability and robustness.
II-B Design of the Error Amplification Module (EAM)
Advancements in technology have significantly improved the sampling rates and resolution of ADCs, making fully digital laser repetition frequency locking systems based on high-speed ADCs feasible[23]. In applications such as optical frequency combs, laser repetition frequencies can reach tens to hundreds of MHz, or even GHz[25]. However, typical fully digital laser locking systems operate with ADC sampling rates around 100 MHz, which, according to the Nyquist theorem, limits their ability to lock higher repetition frequency lasers. While utilizing higher sampling rates and resolutions could improve performance, the high costs make this impractical for many applications. To accommodate lasers with higher repetition frequencies, common techniques involve down-converting using methods such as frequency divider and mixing. Frequency divider lowers the frequency of the error signal during down-conversion, which can lead to precision loss, Mixing, on the other hand, requires signal synthesizers and mixers, which may introduce additional noise.
To improve precision while reducing system costs, an EAM is designed to preprocesses the laser signal by incorporating principles of frequency multiplication and beat frequency methods. The PLL-based CFM multiplies the laser repetition frequency, while a PLL-based CFG, equipped with jitter-cleaning functionality, generates stable reference signals from the atomic clock for mixing and provides clock signals to the DFLM. Both the CFG and CFM feature zero-delay functionality, reducing system latency and increasing bandwidth.
While some approaches extract higher harmonics from the laser signal for mixing to obtain higher error frequencies, this requires BFP with narrow pass band, leading to poor out-of-band rejection, increased costs, and potential spectral aliasing, which may compromise measurement precision. Additionally, as the order of laser harmonics increases, the signal amplitude diminishes, and using high-gain operational amplifiers to boost the signal can introduce additional noise, making high-precision measurements more challenging.
By amplifying and down-converting the laser repetition frequency error using the EAM, measurement precision is enhanced while system costs are reduced, making the method more practical.
The input laser repetition frequency signal undergoes band-pass filtering to isolate the desired repetition signal:
${{f}_{x}}={{f}_{0}}+\Delta f$ (1)
Where ${{f}_{0}}$ is the expected laser repetition frequency, and $\Delta f$ is the frequency difference between the expected and actual laser repetition frequency.
The CFM multiplies the laser repetition frequency by a factor of $m$, generating a high-frequency signal $m{{f}_{x}}$.
Simultaneously, the CFG processes the stable reference signal ${{f}_{r}}$ from the atomic clock to produce a stable mixing signal:
${{f}_{m}}=m{{f}_{0}}-{{f}_{r}}$ (2)
The high frequency input signal $m{{f}_{x}}$ and the stable mixing signal ${{f}_{m}}$ are then mixed and passed through an LPF, producing an amplified error signal:
${{f}_{e}}={{f}_{r}}+m\Delta f$ (3)
This amplified error signal, which matches the reference frequency, reduces the impact of back-end noise, thereby improving measurement precision.
The EAM can be configured to accommodate different laser repetition frequencies by adjusting the multiplication factor and the frequency of the mixing signal generated by the PLL, ensuring the processed signal can be effectively sampled by the ADC.
Figure 2: Block diagram of measurement methods. (a) DQD method (b) DMTD method. REF: reference source.
II-C Design of DMTD and Frequency Extraction Module
In traditional laser repetition frequency locking systems, the digital quadrature demodulation (DQD) method, shown in Fig.2(a), is widely used for frequency measurement due to its simple structure and high resolution[20]. However, it is susceptible to inaccuracies caused by the NCO frequency generator. In contrast, DMTD method [26] can counteract noise from the common reference source, offering higher accuracy and flexible sampling period[27].
The analog DMTD method measures the phase difference of beat signals using zero-crossing detection and time interval counters[28][29]. This approach is prone to noise interference at the zero-crossing points, leading to measurement errors. Additionally, the presence of numerous analog components introduces extra noise into the system. Traditional DMTD methods control the measurement period by adjusting the beat frequency[30], which can reduce measurement resolution when the period is shortened.
This paper employs an ADC-based digital DMTD method Fig.2(b) to measure the frequency difference between two signals. This method avoids the zero-crossing noise issue associated with analog methods[31] and utilizes digital components such as mixers and filters. Digital components offer high consistency. In the symmetric design of the DMTD method, this consistency maximizes the cancellation of common-mode noise, significantly reducing noise and enhancing measurement accuracy. What’s more, this method does not require the two signals to have identical nominal frequencies, allowing for the measurement of signals with significant frequency deviations in the MHz range, which enhances the system’s measurement tolerance.
The process begins by sampling both the amplified error signal and the reference signal using an ADC. The amplified error frequency signal is represented as:
${{U}_{e}}\left(t\right)={{A}_{e}}sin\left[2\pi{{f}_{e}}t+{{\varphi}_{e}}+\Delta{{\varphi}_{e}}(t)\right]$ (4)
while the reference signal from a highly stable source is expressed as:
${{U}_{r}}\left(t\right)={{A}_{r}}sin[2\pi{{f}_{r}}t+{{\varphi}_{r}}+\Delta{{\varphi}_{r}}(t)]$ (5)
Here, ${{A}_{e}}$ and ${{A}_{r}}$ denote the amplitudes, ${{f}_{e}}$ and ${{f}_{r}}$ represent the frequencies, and ${{\varphi}_{e}}$, ${{\varphi}_{r}}$ are the initial phases. The terms $\Delta{{\varphi}_{e}}(t)$ and $\Delta{{\varphi}_{r}}(t)$ indicate the random phase noise in the signals.
Two orthogonal signals are then generated using the NCO within the FPGA:
${{U}_{s1}}\left(t\right)={{A}_{s1}}sin\left(2\pi{{f}_{s1}}t+{{\varphi}_{s1}}+\Delta{{\varphi}_{s1}}(t)\right)$ (6)
${{U}_{s2}}\left(t\right)={{A}_{s2}}cos\left(2\pi{{f}_{s2}}t+{{\varphi}_{s2}}+\Delta{{\varphi}_{s2}}(t)\right)$ (7)
Since the two orthogonal signals are generated by the same NCO, ideally, they satisfy ${{A}_{s1}}={{A}_{s2}}$, ${{f}_{s1}}={{f}_{s2}}={{f}_{NCO}}$, and ${{\varphi}_{s1}}={{\varphi}_{s2}}$. The random phase jitter contained in the orthogonal signals, represented by $\Delta{{\varphi}_{s1}}\left(t\right)$ and $\Delta{{\varphi}_{s2}}\left(t\right)$, is also symmetric and satisfies $\Delta{{\varphi}_{s1}}\left(t\right)=\Delta{{\varphi}_{s2}}\left(t\right)$ under ideal conditions.
Next, the reference signal ${{U}_{r}}$ is digitally mixed with the NCO signals and low-pass filtered, yielding digital signals with low-frequency components:
${{I}_{1}}\left(t\right)\approx{{A}_{I1}}\sin\left[2\pi{{f}_{h1}}t-{{\varphi}_{s2}}+{{\varphi}_{r}}-\Delta{{\varphi}_{s2}}(t)+\Delta{{\varphi}_{r}}(t)\right]$
${{Q}_{1}}\left(t\right)\approx{{A}_{Q1}}\cos\left[2\pi{{f}_{h1}}t-{{\varphi}_{s1}}+{{\varphi}_{r}}-\Delta{{\varphi}_{s1}}(t)+\Delta{{\varphi}_{r}}(t)\right]$
Here, ${{f}_{h1}}={{f}_{r}}-{{f}_{s1,2}}$. The same operations are applied to the amplified error frequency signal ${{U}_{e}}$:
${{I}_{2}}\left(t\right)\approx{{A}_{I2}}\sin\left[2\pi{{f}_{h2}}t-{{\varphi}_{s2}}+{{\varphi}_{e}}-\Delta{{\varphi}_{s2}}(t)+\Delta{{\varphi}_{e}}(t)\right]$
${{Q}_{2}}\left(t\right)\approx{{A}_{Q2}}\cos\left[2\pi{{f}_{h2}}t-{{\varphi}_{s1}}+{{\varphi}_{e}}-\Delta{{\varphi}_{s1}}(t)+\Delta{{\varphi}_{e}}(t)\right]$
where ${{f}_{h2}}={{f}_{e}}-{{f}_{s1,2}}$.
The phase difference is derived from demodulating IQ signals as:
${{\theta}_{1}}\left(t\right)$ $={{\tan}^{-1}}\frac{{{I}_{1}}\left(t\right)}{{{Q}_{1}}\left(t\right)}$ (12)
$=2\pi{{f}_{h1}}t-{{\varphi}_{s1}}+{{\varphi}_{r}}-\Delta{{\varphi}_{s1}}+\Delta{{\varphi}_{r}}$
${{\theta}_{2}}\left(t\right)$ $={{\tan}^{-1}}\frac{{{I}_{2}}\left(t\right)}{{{Q}_{2}}\left(t\right)}$ (13)
$=2\pi{{f}_{h2}}t-{{\varphi}_{s2}}+{{\varphi}_{e}}-\Delta{{\varphi}_{s1}}+\Delta{{\varphi}_{e}}$
Where ${{\theta}_{1}}\left(t\right)$ is the phase difference between the reference signal ${{U}_{r}}$ and the two orthogonal signals generated by the NCO. Similarly, ${{\theta}_{2}}\left(t\right)$ represents the phase difference between the amplified error frequency signal ${{U}_{e}}$ and the two orthogonal signals generated by the NCO.
By subtracting ${{\theta}_{2}}(t)$ from ${{\theta}_{1}}(t)$, the phase difference between the amplified error frequency signal and the reference signal is calculated as:
$\theta\left(t\right)$ $={{\theta}_{1}}(t)-{{\theta}_{2}}(t)$ (14)
$=2\pi\left({{f}_{r}}-{{f}_{e}}\right)t+{{\varphi}_{r}}-{{\varphi}_{e}}+\Delta{{\varphi}_{r}}\left(t\right)-\Delta{{\varphi}_{e}}\left(t\right)$
The frequency difference between the reference signal and the amplified error signal can be calculated by finding the derivative of phase difference with time $t$:
${{f}_{r,e}}$ $=\frac{\theta{{\left(t\right)}^{\prime}}}{2\pi}={{f}_{r}}-{{f}_{e}}+\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}$ (15)
$=m\Delta f+\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}$
This frequency difference includes high-frequency phase noise $\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}$, which introduces measurement errors. However, the frequency difference $m\Delta f$ remains relatively stable over short periods, while the phase noise decreases when averaged. Thus, by averaging over short periods, the phase noise decreases, and the frequency difference stabilizes:
${\overline{{f}_{r,e}}}=m\Delta f+\overline{\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}}$ (16)
With sufficient averaging, the phase noise diminishes, and the frequency difference approaches $m$, providing a stable input for digital feedback algorithms that lock the laser repetition frequency.
Figure 3: Phase unwrapping diagram. CNT: counter.
Due to inevitable differences between the NCO signal, the measured signal, and the reference signal, the phase between them slowly but steadily drifts, causing phase wraps. During frequency calculation, phase unwrapping is required to obtain accurate data. Differentiating the phase data can convert it into instantaneous frequency. Then averaging the frequency difference can reduce the impact of phase noise.
However, random-access memory (RAM) resources in the FPGA are limited, direct averaging over long time periods (e.g., 100 million samples per second at a 100 MHz ADC sampling rate) is impractical.
Additionally, different feedback devices operate at varying response speeds, and it may be necessary to adjust the measurement period to meet specific requirements. To address these challenges and minimize phase noise, a counter-based phase unwrapping method is employed for frequency calculation and decimation. This method requires only a counter and minimal RAM, allowing for frequency averaging over a wide range. Fig.3 illustrates the phase unwrapping process, where the counter increments or decrements depending on the direction of the phase wrap. By adding the counter’s phase value to the current phase and differentiating over different time intervals, frequency measurements can be obtained.
In a digital system, signals consist of discrete sampled points. Based on the derived approach above, a series of phase points is obtained within the FPGA, denoted as ${\theta(1),\theta(2),\ldots,\theta(n-1),\theta(n)}$. The frequency at any given point is expressed as:
$f(n)=\frac{\theta(n)-\theta(n-1)}{2\pi}.$ (17)
Averaging the frequency over multiple points results in:
$\overline{f}=\frac{1}{2\pi m}\sum_{n=2}^{m}\left(\theta(n)-\theta(n-1)\right),$ (18)
where $m$ is the length of the measurement period. Simplifying this expression yields:
$\overline{f}=\frac{\theta(m)-\theta(1)}{2\pi m}.$ (19)
Thus, only the first and last phase values need to be stored in the FPGA to calculate the frequency over the cycle of $m$. Using a counter, the unwrapped phase values $\theta^{\prime}(1)$ and $\theta^{\prime}(m)$ can be determined as follows:
$\theta^{\prime}(1)=\theta(1),$ (20)
$\theta^{\prime}(m)=\theta(m)+2\pi\text{CNT}(m).$ (21)
Finally, the average frequency difference is expressed as:
$\overline{f}=\frac{\theta^{\prime}(m)-\theta^{\prime}(1)}{2\pi m}.$ (22)
This counter-based phase unwrapping method allows for frequency measurements over periods ranging from nanoseconds to seconds, with adjustable feedback periods to suit different system requirements.
II-D Feedback Loop
In the feedback loop, the amplified frequency difference serves as the error signal for laser stabilization. It is first processed by an LPF to remove high-frequency noise. A PID controller then computes the difference between the error signal and the settled frequency offset. This enables flexible locking of the laser to different relative frequencies. The calculated error is converted to an output voltage by the DAC, and the voltage is amplified by a high-voltage amplifier before being applied to the piezoelectric transducer (PZT) to adjust the laser’s frequency. As all feedback modules are digital, the system can be easily configured to support different feedback devices such as the PZTs and the servos.
III Implementation and Performance
III-A Implementation
Figure 4: Block diagram of the error amplification module.
Figure 5: Function modules of FPGA logic firmware. NCO: numerically controlled oscillator.
Figure 6: Annotated photograph of the high-speed digital FPGA board.
To evaluate the practicability and the performance of the proposed method, a laser repetition frequency locking system was designed, including the EAM and a high-speed digital FPGA board. Fig.4 shows the block diagram of the EAM and the function modules of the FPGA logic firmware are shown in Fig.5. The laser used in the experiment has a repetition frequency of 50 MHz. Considering that the CFG and CFM used in the system can operate below 2 GHz, a 20x frequency multiplication was applied in the EAM. This effectively amplifies the frequency error by a factor of 20 for subsequent measurements. Fig.6 shows the annotated photograph of the high-speed digital FPGA board. A Xilinx Kintex-7 FPGA is used as the central control unit, with the platform’s digital signal control entirely managed by the programmable digital logic within the FPGA. The power supply section includes both analog and digital circuits to generate the main power supply. Communication between the computer and the FPGA is facilitated through a 1G Ethernet interface. The PLL based CFG provides clock signals to the FPGA, generates ADC sampling signals, and supplies the local oscillator signals for external analog mixing. Additionally, the system includes two CFMs to support multiple channel frequency stabilization. The platform integrates a four-channel high-speed ADC with a resolution of 16 bits and a sampling rate of 100 MS/s. A high-precision DAC, connected via an expansion interface, features a resolution of 20 bits and an output frequency of 1 MS/s. Multiple external expansion interfaces are available for connecting additional DACs and other devices such as DDS. All input and output channels are equipped with $50-\Omega$ Sub-Miniature Version A (SMA) terminations. A 4-GB DDR3 memory provides sufficient data storage capacity and memory bandwidth.
The DMTD frequency extraction method as well as the feedback is deployed in the FPGA for frequency locking. What’s more, a Lock status Detector module can detect the frequency difference in real-time, outputting a locked signal when the frequency error remains within a specified range for a certain period. The Ethernet Interface allows configuration of the EAM and DFLM by computer. Configurable parameters include the multiplication factor, CFG output frequency, frequency measurement period in the Frequency Extraction module, the frequency offset between the laser and reference, low-pass filter parameters, PID parameters, and DAC feedback speed. The EAM is adaptable to different laser repetition frequencies, and the backend feedback module can be configured to accommodate various feedback devices such as PZTs and servos.
This system is suitable for locking a single laser and can be expanded to lock dual laser combs by adding more DACs. The reprogrammable FPGA chip offers excellent flexibility and scalability to meet various requirements. Furthermore, the proposed method offers low latency and high precision, making it suitable for various implementations. By integrating higher-precision ADCs and faster DACs, the system’s performance can be further improved. Additionally, coordinating the use of high-speed and low-speed DACs may enhance locking performance.
III-B DFLM Noise Floor
Figure 7: Schematic diagram of DFLM noise floor measurement.
Fig.7 illustrates the setup for measuring the DFLM noise floor. A reference source is split using a power divider and fed into both channel of the DFLM module through equal-length RF cables. The data is then transmitted via Ethernet Interface to the local PC for stability analysis. Since the two input sine wave signals originate from the same source, their phase difference remains at zero, making this configuration ideal for measuring the system’s noise floor. The reference source used is a rubidium atomic clock, the SAFRAN LPFRS-01, with a nominal frequency of 10 MHz and a stability of $2\times 10^{-11}$ at 1 second. The sampling interval was set to 100 kHz, with a total measurement duration of 400 seconds.
Figure 8: Time deviation in DMTD method and DQD method.
The precision of the DMTD method described in this paper was compared with the DQD method. Both methods use the same external test circuit structure, but the DMTD frequency extraction module is replaced with the DQD frequency extraction module in DQD noise floor test. Fig.8 presents the measured DFLM noise floor, displaying the results both before and after filtering for both methods. The figure shows the unfiltered noise floor and the noise floor after applying a 100 Hz low-pass filter for both methods. The DMTD method demonstrated better noise floor both before and after filtering. After filtering, the DMTD method achieved a noise floor of approximately $1.58\times 10^{-13}@1s$, while the traditional DQD method only reached $9.16\times 10^{-12}@1s$. Additionally, the noise floor of the phase noise analyzer (Microchip model 53100A) are also shown in the figure, taken at a 1 kHz measurement period with a 500 Hz low-pass filter. The analyzer’s noise floor was $7.96\times 10^{-14}$ at a gate time of 1s.
The results demonstrate that the DFLM using the DMTD method achieves a noise floor comparable to that of a commercial phase noise analyzer, offering more precise frequency measurements.
III-C Laser Stability Experiment
Figure 9: Laser repetition frequency offset and Allan deviation before and after locking.
Figure 10: Scene of the laser frequency locking experiment.
In a standard laboratory environment, a laser frequency locking experiment was performed using a custom-built mode-locked femtosecond fiber laser. Fig.10 shows the scene of the laser frequency locking experiment. The laser operates at a center wavelength of 1550 nm, with a repetition rate of 50 MHz and a tunable range of ±400 Hz, delivering an average optical power of 100 mW. Two 4 mW optical signals were extracted via a fiber coupler and connected to PDs for optical-to-electrical conversion. One signal was directed to the system for frequency locking, while the other was sent to a commercial phase noise analyzer to measure the frequency difference.
The laser repetition frequency was compared to a rubidium atomic clock to generate the error signal. A 500 Hz LPF was applied to the error signal to remove high-frequency noise, and the filtered signal was then sent through a DAC to control the laser repetition frequency by adjusting the PZT. The feedback frequency was set to 1 kHz, which was limited by the PZT.
To evaluate the frequency locking performance and long-term stability, a Microchip phase noise analyzer (model 53100A) was used to measure the frequency difference. Both the reference for frequency locking and the analyzer were connected to a high-stability rubidium atomic clock with a nominal stability of $2\times 10^{-11}@1s$. The gate time was initially set to 0.001 seconds due to the large free-running frequency difference, and later adjusted to 1 second after locking.
Data were collected for two hours both before and after locking. The results are displayed in Fig.9. The left figure shows the laser’s frequency offset before and after locking, demonstrating a significant reduction in frequency fluctuations once locked. On the right, the figure presents the Allan deviation, where the laser’s short-term Allan deviation at 1 second improved from $6.05\times 10^{-8}$ before locking to $1.04\times 10^{-11}$ after. For long-term stability, the laser achieved a stability of $1.12\times 10^{-12}$ at an averaging time of 10 seconds after locking, a substantial improvement from the before-locking stability of $1.51\times 10^{-7}$. Additionally, the right figure shows the noise floor of the phase noise analyzer used in the measurement, as well as results obtained by measuring two different atomic clocks with the same nominal stability. The locked laser exhibited short-term stability comparable to that of the atomic clock, while its long-term stability exceeded that of the atomic clock.
III-D VCO Stability Experiment
Due to the limited stability as well as the feedback bandwidth of our custom-built laser in the Laser Stability experiment, we were unable to fully demonstrate the system’s precision and performance. To address this, the VCO Stability experiment was introduced. In this setup, a VCO was used to simulate the laser’s behavior, as it shares similar characteristics—namely, a stable output frequency and voltage-controlled adjustability. The VCO offers higher feedback bandwidth and stability compared to our laser, making it an effective benchmark for evaluating the system’s performance under current experimental conditions.
The VCO used in this test is the AOCJY1 model from Abracon LLC, with a nominal frequency of 10 MHz, a frequency adjustment range of ±100 Hz, and a nominal stability of $1\times 10^{-9}@1s$. The system was locked using the rubidium atomic clock mentioned earlier as the high stability reference source, which has a nominal stability of $2\times 10^{-11}@1s$. A 300 Hz low-pass filter was employed to filter the error signal, eliminating high-frequency noise, while the feedback frequency was set to 10 kHz.
Figure 11: Fig(a) illustrates the frequency offset of the VCO relative to the reference frequency over 8000 seconds, showing results both before and after locking, as well as the frequency control portion. Fig(b) offers a zoomed-in view of the frequency fluctuations after locking, with a standard deviation measured at 0.0077 mHz. Fig(c) shows the Allan variance before and after locking, with the green line representing the before-locking state and the red line depicting the after-locking state.
The same phase noise analyzer (model 53100A) was used to measure the frequency difference, with the gate time set to 1 second. The rubidium atomic clock was again used as the reference source for the system and the phase noise analyzer. The total measurement time was 6 hours, with 8000 seconds selected for display. As shown in Fig.11(a), from 0 to 836 seconds, the VCO operated in free-run mode. At 837 seconds, the FPGA initiated frequency locking, successfully locking the VCO to the reference frequency within 1 second, maintaining the lock for the remainder of the test. At the end of the figure, the system’s frequency control capability is demonstrated by switching between two different frequency offsets, achieving real-time and precise frequency locking. Unlike traditional phase-locked loops, the system allows for arbitrary frequency locking within the allowable precision range, offering excellent frequency stability with minimal parameter adjustment required.
Fig.11(b) provides a zoomed-in view of the frequency fluctuations after locking, with a standard deviation of 0.0077 mHz. Fig.11(c) presents the Allan deviation before and after locking, with the green line indicating the before-locking state and the red line showing the after-locking state. The short-term Allan deviation at 1 second for the locked VCO frequency reached $9.48\times 10^{-13}$, a significant improvement over the $2.73\times 10^{-10}$ observed before locking. In terms of long-term stability, the VCO exhibited a stability of $1.19\times 10^{-9}$ at an averaging time of 10 seconds before locking, which improved to $9.58\times 10^{-14}$ after locking. The locked VCO demonstrated exceptional stability, confirming the system’s accuracy and reliability.
IV Conclusion
In this paper, we have presented an FPGA-based versatile and digitalized method for pulse laser repetition frequency locking. This method can realize precise frequency measurement, providing reliable frequency lock with minor parameter adjustments using DMTD techniques. The FPGA-based architecture as well as the EAM enables flexibility and adaptability, allowing the system to be reconfigured for various feedback devices and laser frequencies. We design a high-speed FPGA board to evaluate the practicability and the performance of the method.
The results from the Laser Stability experiment demonstrate the system’s capability to significantly reduce frequency fluctuations, achieving a long-term Allan deviation improvement from $1.51\times 10^{-7}$ to $1.12\times 10^{-12}$ at 10 second. The VCO Stability experiment further confirms the system’s precision, with a long-term Allan deviation of $9.58\times 10^{-14}$ at 10 second after locking, making it highly suitable for applications requiring long-term stability.
All these demonstrate that the proposed laser repetition frequency locking method offers a precise, reconfigurable, and compact solution for laser frequency stabilization, with significant potential for future enhancements in both hardware and software.
Acknowledgments
The authors would like to thank Qing Sun at National Institute of Metrology for useful discussions.